OpenStudy (anonymous):

divided, if the remainder is or is equal to zero the it shows that the binomial is a factor of the polynomial can anyone show me an example of this please

OpenStudy (anonymous):

pleeeeeaaaaaaassssseeeeee

OpenStudy (anonymous):

Divide \[8x ^{3}-1/ (2x-1) = 4x ^{2}+2x+1\] In General \[f(x)/g//x)=q(x)+r(x)/d(x)\] In example above r(x)=0